# Strategic Learning and Trading in Broker-Mediated Markets

### v1

@@ 5.1   Performance and comparison against market practices
    Benchmark 1: externalise the trading flow of the informed trader, internalise the trading flow of the uninformed trader, and unwind the broker’s inventory with TWAP. For this benchmark, the trading speed of the broker is:
@@
Below, we use $Q_{t}^{B,(i)}$, for $i\in\{1,2,3\}$, to denote the inventory of the broker when she follows strategy $\nu_{t}^{(i)}$.
@@
1 0 (570) $>0.5$
2 38 (354) $<0.001$
3 95 (695) $<0.001$
Table 1: Model outperformance with respect to the three benchmarks and their $p$-values from a one sided $t$-test ($H_{0}:\,\text{Out}^{(i)}=0$ and $H_{1}:\,\text{Out}^{(i)}>0$). The raw performance of the optimal strategy is 0.32 (2.81), and the performance for three benchmarks are 0.32 (2.39), 0.19 (2.71), and -0.04 (0.21).
Due to the low signal-to-noise ratio, the performance of the optimal strategy is indistinguishable from that of the first benchmark strategy which externalises the informed flow, internalises the uninformed flow, and unwinds inventory linearly. Note that we do not reject the null hypothesis (that the means are the same) in the $t$-test. In the next section, we study the performance of the broker’s strategy when she employs other sources of information to build her estimate of the private signal.
5.2   Performance with alternative estimators
Motivated by the poor performance of the broker when she estimates the signal from prices, we study the performance of the adaptive strategy (4.77) when the broker employs other estimates of the signal, including that described in Section 4.1.2.
5.2.1 Filtering directly from the informed trader’s speed
Here, we show that when the broker employs the estimate (4.28), which is constructed from the trading flow of the informed trader, she improves her trading performance significantly against the benchmarks described above. [^10] The performance reported here assumes no mispecification of inventory. Table 2 reports the outperformance. The performance we report here is a lower bound to the performance of the optimal strategy which uses the estimate (4.28) because the strategy we study here would be admissible in the adaptive control problem.
@@ 5.2.1 Filtering directly from the informed trader’s speed
1 36 (117) $<0.001$
2 78 (389) $<0.001$
3 127 (518) $<0.001$
Table 2: Alternative estimation of $\alpha$ filtering directly from the informed trader’s speed. Model outperformance with respect to the three benchmarks and their $p$-values from a one sided $t$-test ($H_{0}:\,\text{Out}^{(i)}=0$ and $H_{1}:\,\text{Out}^{(i)}>0$).
We remark that the estimate (4.28) is only valid if (i) the informed trader carries out his trading exclusively with the broker and (ii) that he does not hide information or mislead the broker. If this is not the case, then the inventory of the informed trader, from the perspective of the broker, is mispecified, and the calculation of $\hat{\alpha}^{\text{alt}}_{t}$ above blows up at the terminal time $T$ as the limit in (4.16) blows up due to unaccounted inventory differences.
Figure 5: Left panel shows one sample path for the signal $\alpha$ and the new estimates $\hat{\alpha}$. The right panel shows the misspecified case for the informed trader’s inventory.
The left panel of Figure 5 shows the estimate assuming no mispecification in inventory, and the right panel shows a sample path when $Q^{I}_{0}\sim\mathcal{N}(0,1)$ and the broker assumes that $Q^{I}_{0}=0$. The mispecification of inventory has less impact on the accuracy of the estimate $\hat{\alpha}^{\text{alt}}$ in the beginning of the trading window because the informed trader trades less aggressively at this stage. However, as trading progresses, the accuracy deteriorates. Therefore, if the broker employs $\hat{\alpha}^{\text{alt}}$, then she must rely less on the signal less as the trading nears its close, since inventory mispecification increasingly contaminates the extracted signal at that stage.
Figure 6 shows the four components in the broker’s speed when she employs the new estimates $\hat{\alpha}^{\text{alt}}$. We observe that all components now contribute significantly to her trading strategy. In particular, the speculative component from exploiting $\hat{\alpha}^{\text{alt}}$ is more predominant than in Figure 4.
Figure 6: The four additive components of the broker’s strategy if she employs the new filtering strategy.
@@ 5.2.2 Naive estimate for the signal assuming near-zero values for $\hat{\nu}$
In practice, given that the permanent impact parameter $\mathfrak{p}$ of the broker’s trading activity in the market is small, the noise driving price movements obscures the filtered order flow of the broker. Thus, the contributions of the filtered trading speed $\hat{\nu}$ in the informed trader’s strategy are small enough so that they may be assumed to be zero. We illustrate this in Figure 7.
@@
The broker could obtain a better, near-perfect, estimate for $\alpha$. As shown in Figure 7, the contribution of $\hat{\nu}$ in the informed trader’s strategy is near-zero. Consequently, the broker could ignore those contributions (by assuming $f_{2}(t)\,\hat{\nu}_{t}=0$) and obtain the following alternative estimates for $\alpha$
@@
from L’Hôpital’s rule and the differential equations (3.20)–(3.21). Similar to the filter $\hat{\alpha}^{\text{alt}}$, we remark that this estimate is only valid if the informed trader carries out his trading exclusively with the broker. If this is not the case, then the inventory of the informed trader, from the perspective of the broker, is mispecified, and the calculation of $\hat{\alpha}_{t}$ above blows up at the terminal time $T$.
Figure 8 shows the differences between $\hat{\alpha}^{\textup{alt}}_{t}$ and $\hat{\alpha}^{\textup{naive}}_{t}$ with (right panel) and without (left panel) inventory mispecification. The differences in the left panel are negligible for practical purposes (the $y$-axis is in units of $10^{-5}$), and the differences in the right panel are in the same order of magnitude as the mispecification. As expected, the differences blow up when $t\uparrow T$.
Figure 8: Both panels show the differences between $\hat{\alpha}^{\textup{alt}}_{t}$ and $\hat{\alpha}^{\textup{naive}}_{t}$ with a 90% confidence band over simulations. Left panel is for no mispecification of inventory. The right panel has mispecification of inventory, that is, $Q^{I}_{0}\sim\mathcal{N}(0,1)$ and the broker assumes that $Q^{I}_{0}=0$.
Finally, we compute the outperformance over the benchmarks when the broker employs the estimate $\hat{\alpha}^{\textup{naive}}$ of $\alpha$ above, and assuming no mispecification of inventory. As expected, we find that the outperformances are the same as those in Table 2. That is, if we compare the performance obtained using the two filtered signals in this section, a $t$-test does not find evidence to reject the null hypothesis that their mean performances are the same.
@@ 5.3   Second-order effects
In the problem formulation above, both agents act strategically, though the informed trader’s strategy is only strategic to a first order. Specifically, the informed trader does not account for the possibility that the broker may filter his private signal. Conversely, the broker operates under the assumption that the informed trader will incorporate her trading activity to enhance trading performance. Here, we explore how to incorporate this second-order effect into the informed trader’s problem and evaluate the value of the broker’s second-order effect.
5.3.1 Informed trader
From the perspective of the informed trader, a second-order effect can be incorporated by assuming that the broker’s trading follows the model
$\nu_{t}=c_{t}\,\eta_{t}+\underline{\nu}_{t}\,,$ (5.6)
where $c_{t}$ is a deterministic function that reflects how the broker externalises his trading flow, and $\underline{\nu}_{t}$ is the remainder of the broker’s trading flow. In this setup, the informed trader assumes that $\underline{\nu}$ is exogenous and mean reverting, as in the initial model. Given this assumption, and similar to the approach above, the informed trader can filter $\underline{\nu_{B}}$ to compute $\hat{\underline{\nu}}$, and then formulate an adaptive control problem. This control problem would be similar to that in (3.10), but where one changes $\hat{\nu}$ to $\underline{\hat{\nu}}$, and adds a term of the form $\mathfrak{p}\,c_{t}\,\eta_{t}\,\partial_{s}H^{I}$. While we omit the detailed calculations here, the intuition is that the informed trader benefits in this modified model, as he effectively “observes” more of the broker’s trading speed, which enhances his strategic advantage.
$\nu_{t}=c_{t}^{1}\,\hat{\alpha}_{t}+c_{t}^{2}\,Q^{I}_{t}+\underline{\nu}_{t}\,,$ (5.7)
$\mathrm{d}S_{t}=\left[\mathfrak{p}\,\hat{\underline{\nu}}_{t}+(1+\mathfrak{p}\,c_{t}^{1})\,\alpha_{t}+\mathfrak{p}\,c_{t}^{2}\,Q^{I}_{t}\right]\,\mathrm{d}t+\sigma^{S}\,\mathrm{d}\tilde{W}^{S,I}_{t}\,,$ (5.8)
where $\hat{\underline{\nu}}_{t}$ is the filter of ${\underline{\nu}}_{t}$.
5.3.2 Broker
$\eta^{*}_{t}=f_{1}(t)\,\hat{\alpha}_{t}+c\,f_{2}(t)\,\nu_{t}+f_{3}(t)\,Q^{I*}_{t}\,,$ (5.9)
Based on the model assumption (5.9), the broker computes her optimal speed $\nu^{*,c}$. Note that the case $c=1$ corresponds to the initial model described in Section 4. Our aim is now to examine how $\nu^{*,c}$ deviates from the optimal strategy in the original model where $c=1$. To study the manipulation effects, we reduce the influence of the risk terms on the strategies of both the broker and the informed trader by setting $\beta_{0}^{I}=\beta_{0}^{B}=10^{-5}$.
The broker’s trading speed is a component in the informed trader’s strategy (5.9). Thus, in our original model, the assumption (5.9) may be interpreted as a certain “belief” of the broker in her ability to influence the informed trader’s strategy. Figure 9 shows that the broker uses this component to her advantage. More precisely, the broker will, when profitable, intentionally mislead the informed trader into trading in the opposite direction of her inventory. The aim is to offload her inventory to her client rather than unwinding it in the lit market herself.
@@ 5.4   Mispecification of learning parameters
Here, we examine how mispecifications in model parameter values affect the outperformance metrics reported above. Using the values reported in Table 1 as a baseline, we stress each parameter by increasing and decreasing its value by $50\%$, and report the results in Table 3.
@@
1 0 (570) -11 (569) -12 (573) -12 (573) -11 (566)
2 38 (354)* 38 (356)* 37 (362)* 37 (362)* 38 (351)*
3 95 (695)* 88 (693)* 87 (696)* 87 (696)* 88 (691)*
Table 3: Outperformance of model to the three benchmarks when the broker mispecified her learning parameters. Significant cases are marked with *.
We observe that, even with parameter mispecification, the broker’s optimal strategy continues to outperform benchmarks 2 and 3, with outperformance levels remaining approximately the same as in the baseline. However, the broker’s outperformance over the first benchmark now turns negative; note that these values are not statistically significant.
@@ 5.5   The role of externalisation and stressing model parameters
Remark 3.
If the broker is correct in her estimate of $\alpha$ and if the informed trader only conducts his trade with the broker (the broker has a good estimate of $Q^{I}$), then, the broker’s “externalisation” that happens in advance of the trades taking place, might be catalogued as “frontrunning”. For example, the broker might buy (resp. sell) in the lit market in advance of the informed trader sending his buy (resp. sell) instruction. However, we remark that these actions are taken based on a probabilistic model as opposed to being taken based on factual knowledge of what the client is going to do next.
@@
Figure 10 shows that the mean-reversion parameter $\kappa^{\alpha}$ in the signal dynamics influences the effective externalisation rate significantly. Large values of $\kappa^{\alpha}$ result in a signal that diffuses less. As such, the broker trusts her estimates more and increases her speculation part, thus she externalises more. As the signal diffuses less (from zero), its role diminishes in the strategies of both the broker and informed trader; see Figure 11. Visually, the signal appears to impact both the broker and informed trader’s strategies similarly. However, the broker’s sensitivity to the signal decreases at a faster rate when $\kappa^{\alpha}$ is low (i.e. when the signal diffuses more). This can be attributed to the broker’s growing uncertainty in her estimate of the signal due to accumulated errors. When $\kappa^{\alpha}$ is high (i.e. when the signal diffuses less), the broker’s sensitivity to the signal decreases at a more equal rate to the trader. Consequently, the broker increases her externalisation rate without concern for the risks associated with signal misspecification.

### v2

@@ 5.1  Performance and comparison against benchmarks
    Benchmark 1: externalise the trading flow of the informed trader, internalise the trading flow of the uninformed trader, and unwind remaining inventory with TWAP. For this benchmark, the trading speed of the broker is:
@@
1 0 (570) >0.5>0.5
2 38 (354) <0.001<0.001
3 95 (696) <0.001<0.001
@@
1 36 (117) <0.001<0.001
2 78 (389) <0.001<0.001
3 127 (518) <0.001<0.001
Table 1: Model outperformance with respect to the three benchmarks and their $p$-values from a one sided $t$-test ($H_{0}:\,\text{Out}^{(i)}=0$ and $H_{1}:\,\text{Out}^{(i)}>0$). The left table is the first strategy when the broker filters from prices; the right table is when she employs the alternative filter. The raw performance of the optimal strategy employing the first learning method is 0.32 (2.81). For the alternative learning equation, the performance is 0.45 (2.08). The performance for three benchmarks are 0.32 (2.39), 0.19 (2.71), and -0.04 (0.21).
Due to the low signal-to-noise ratio, the performance of the first strategy is indistinguishable from that of the first benchmark strategy which externalises the informed flow, internalises the uninformed flow, and unwinds inventory linearly. Note that we do not reject the null hypothesis (that the means are the same) in the $t$-test.
On the other hand, the second strategy outperforms all the benchmarks significantly due to better estimates of the signal. We remark that the estimate (4.9) is only valid if (i) the informed trader carries out his trading exclusively with the broker and (ii) that he does not hide information or mislead the broker. If this is not the case, then the inventory of the informed trader, from the perspective of the broker, is mispecified, and the calculation of $\hat{\alpha}^{\text{alt}}_{t}$ above blows up at the terminal time $T$ as the limit in (B.10) blows up due to unaccounted inventory differences.
Figure 4 shows a sample path when $Q^{I}_{0}\sim\mathcal{N}(0,1)$ and the broker assumes that $Q^{I}_{0}=0$. The mispecification of inventory has less impact on the accuracy of the estimate $\hat{\alpha}^{\text{alt}}$ in the beginning of the trading window because the informed trader trades less aggressively at this stage. However, as trading progresses, the accuracy deteriorates. Therefore, if the broker employs $\hat{\alpha}^{\text{alt}}$, then she must rely less on the signal less as the trading nears its close, since inventory mispecification increasingly contaminates the extracted signal the later stages.
Figure 4: One sample path for the signal $\alpha$ and the new estimates $\hat{\alpha}$ when the broker mispecifies the informed trader’s inventory.
@@ 5.1.1 Naive estimate for the signal assuming near-zero values for $\hat{\nu}$
In practice, given that the permanent impact parameter $\mathfrak{p}$ of the broker’s trading activity in the market is small, the noise driving price movements obscures the filtered order flow of the broker. Thus, the contributions of the filtered trading speed $\hat{\nu}$ in the informed trader’s strategy are small enough so that they may be assumed to be zero. We illustrate this in Figure 5 (see middle panel).
@@
The broker could obtain a better, near-perfect, estimate for $\alpha$ given that the contribution of $\hat{\nu}$ in the informed trader’s strategy is near-zero. More precisely, the broker could ignore those contributions (by assuming $f_{2}(t)\,\hat{\nu}_{t}=0$) and obtain the following alternative estimates for $\alpha$
@@
using (A.12)–(A.13). Similar to the filter $\hat{\alpha}^{\text{alt}}$, we remark that this estimate is only valid if the informed trader carries out his trading exclusively with the broker. If this is not the case, then the inventory of the informed trader, from the perspective of the broker, is mispecified, and the calculation of $\hat{\alpha}_{t}$ above blows up at the terminal time $T$.
@@ 5.2  Second-order effects
In the problem formulation above, both agents act strategically, though the informed trader’s strategy is only strategic to a first order. Specifically, the informed trader does not account for the possibility that the broker may filter his private signal. Conversely, the broker operates under the assumption that the informed trader will incorporate her trading activity to enhance trading performance. In B.8 we investigate how to adapt our framework to accommodate these second-order effects.
@@ 5.3  Mispecification of learning parameters
Next, we examine how mispecification of the value of model parameters affect the outperformance metrics reported above. Using the values reported in Table 1 as a baseline, we stress each parameter by increasing and decreasing its value by $50\%$, and report the results in Table 2 when the broker filters from prices and Table 3 when she filters from the informed trader’s flows.
@@
1 0 (570) 0 (569) 0 (573) -1 (573) 0 (566)
2 38 (354)* 39 (352)* 38 (358)* 38 (358)* 39 (348)*
3 95 (696)* 95 (695)* 95 (697)* 95 (697)* 96 (693)*
Table 2: Outperformance of model to the three benchmarks when the broker mispecified her learning parameters for the first strategy. Significant cases are marked with *.
1 36 (117)* 33 (153)* 36 (181)* 35 (200)* 15 (209)*
2 78 (388)* 74 (512)* 78 (315)* 77 (282)* 53 (627)*
3 127 (518)* 124 (529)* 126 (528)* 126 (532)* 108 (549)*
1 36 (117)* 36 (119)* 37 (113)* 37 (112)* 36 (122)*
2 78 (388)* 78 (383)* 78 (404)* 78 (408)* 77 (376)*
3 127 (518)* 126 (518)* 127 (518)* 127 (518)* 126 (518)*
Table 3: Outperformance of model to the three benchmarks when the broker mispecified her learning parameters for the alternative strategy. Significant cases are marked with *.
We observe that, even when the broker mispecifies the parameters when filtering from prices, her outperformance levels remain approximately the same as in the baseline. Thus, benchmark 1 remains a strong candidate in this scenario. However, when the broker filters from the informed trader’s flows, the optimal strategy outperforms all the benchmark, even with parameter mispecification. Thus, filtering from informed trader’s flows remains a strong strategy for the broker.
@@ 5.4  The role of externalisation and stressing model parameters
Figure 6 shows that the mean-reversion parameter $\kappa^{\alpha}$ in the signal dynamics influences the effective externalisation rate significantly. Large values of $\kappa^{\alpha}$ result in a signal that diffuses less. As such, the broker trusts her estimates more and increases her speculation part, thus she externalises more. As the signal diffuses less (from zero), its role in the strategies of both the broker and informed trader diminishes; see Figure 7. Visually, the signal appears to impact both the broker and informed trader’s strategies similarly. However, the broker’s sensitivity to the signal decreases at a faster rate when $\kappa^{\alpha}$ is low (i.e. when the signal diffuses more). This can be attributed to the broker’s growing uncertainty in her estimate of the signal due to accumulated errors. When $\kappa^{\alpha}$ is high (i.e. when the signal diffuses less), the broker’s sensitivity to the signal decreases at a more equal rate to the trader. Consequently, the broker increases her externalisation rate without concern for the risks associated with signal misspecification.
@@
Furthermore, when the value of $\sigma^{B}$ in the informed trader’s model (3.2) is large (or when $\theta^{B}$ is small), the informed trader believes that the broker will often trade aggressively. In response, the broker increases her externalisation rate. To assess the proportion of the average informed flow that the broker externalises, and following the approach in Bergault and Sánchez-Betancourt (2025), we compute the proxy $G(\nu^{*}_{t})/G(\eta^{*}_{t})$ of the quotient $\nu^{*}_{t}/\eta^{*}_{t}$. The function $G$ is defined as $G(x)=\max(x,\epsilon)\,\mathds{1}(x\geq 0)+\min(x,-\epsilon)\,\mathds{1}(x<0)$, with $\epsilon=0.1$. The function $G$ prevents the ratio $\nu^{*}_{t}/\eta^{*}_{t}$ from becoming undefined or excessively large when $\eta^{*}_{t}$ is close to zero.
Figure 8 illustrates that, for the naive case, the externalisation is ranging between $70\%$ and $90\%$. The ordering of the lines when stressing $\kappa^{\alpha}$ in the right panel of Figure 8 is similar to those in Figure 6.
5.5  Discussion: equilibrium of beliefs
In practice, both the broker and the informed trader will adjust their beliefs with time. The broker believes that the signal is characterised by parameters $\kappa^{\alpha,B}$ and $\sigma^{\alpha,B}$ (not necessarily the true parameters $\kappa^{\alpha}$ and $\sigma^{\alpha})$, similarly, the informed trader models the trading speed of the broker as an OU process with parameters $\theta^{B}$ and $\sigma^{B}$. Similar to Section 4 in Donnelly and Li (2025), one could then look at the equilibrium of beliefs. More precisely, if we $\mathfrak{b}^{B}=(\kappa^{\alpha,B},\sigma^{\alpha,B})$ and $\mathfrak{b}^{I}=(\theta^{B},\sigma^{B})$ be the beliefs of the informed trader and the broker respectively, we could study the performances of each player as they tweak these beliefs in search of an equilibrium point. We envisage that such an exercise would yield equilibrium beliefs $\mathfrak{b}^{B*}$ and $\mathfrak{b}^{I*}$ with the property that (i) the broker’s beliefs are close to the true parameters of the signal and (ii) the informed trader’s beliefs fit well the sample paths of the equilibrium trading speed of the broker. This investigation is an interesting avenue for future research.
6 Conclusions
In this paper we computed the optimal strategies of a broker and an informed trader when both parties filter the information they do not have. Brokers hold a slight advantage over informed traders because informed traders rely solely on prices to filter information, whereas brokers have prices and order flow at their disposal. We find that when brokers filter the stochastic drift $\alpha_{t}$ from prices, the performance of the optimal strategy is indistinguishable from a naive strategy that externalises the order flow from the informed, internalises the uninformed flow, and unwinds inventory according to a TWAP schedule. On the other hand, if brokers use the order flow to filter the stochastic drift, the outperformance over the benchmarks is in the order of magnitude of transaction costs.
Funding:
AA’s research is supported by the Oxford-Man Institute of Quantitative Finance through the EPSRC Centre for Doctoral Training in Mathematics of Random Systems: Analysis, Modelling and Simulation (ESPRC Grant EP/S023925/1).
Disclosure statement
The authors have no conflicts of interest to declare that are relevant to the content of this article.
For the purpose of open access, the authors have applied a CC BY public copyright license to any author accepted manuscript arising from this submission.
Appendix A Derivations for the control problem of the informed trader
Here we collect the calculations associated with the control problem of the informed trader.
A.1  Hamilton–Jacobi–Bellman equation
$0$ $=\partial_{t}H^{I}-(\rho_{0}+\rho_{1}\,\mathbb{V}^{I}_{t})\left(q^{I}\right)^{2}+\sup_{\eta}\big[(\mathfrak{p}\,\hat{\nu}+\alpha)\,\partial_{s}H^{I}-\eta\left(s+\mathfrak{b}\,\eta\right)\,\partial_{x^{I}}H^{I}+\eta\,\partial_{q^{I}}H^{I}\big]$
$\qquad+\mathfrak{p}\,\mathbb{V}^{I}_{t}\,\partial_{s\hat{\nu}}H^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\alpha}H^{I}+\frac{1}{2}\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}\partial_{\hat{\nu}\hat{\nu}}H^{I}$ (A.1)
$\qquad+\sigma^{S}\,\sigma^{\alpha}\,\rho\,\partial_{s\alpha}H^{I}+\frac{\mathfrak{p}\,\sigma^{\alpha}\,\mathbb{V}^{I}_{t}\,\rho}{\sigma^{S}}\,\partial_{\alpha\hat{\nu}}H^{I}\,$, (A.2)
with terminal condition
The supremum in the Hamiltonian in (A.2) is attained at
$\eta^{*}=\frac{\partial_{q^{I}}H^{I}-\partial_{x^{I}}H^{I}\,s}{2\,\partial_{x^{I}}H^{I}\,\mathfrak{b}}\,.$ (A.3)
After substitution, employ the ansatz